OpenStudy (anonymous):

x^(log[3](x^2))+(log[3](x^2))^2-10=1/x^2

OpenStudy (anonymous):

\[x ^{\log 3(x ^{2})}+(\log 3(x^{2}))=1/x ^{2}\] Is that what you're looking for?

OpenStudy (anonymous):

add the ^2-10 to log 3(x^2)

OpenStudy (anonymous):

here 3 is base and remaning is N

OpenStudy (anonymous):

-10 is seperate